Gabriele enterprises has bonds on the market making annual payments, with 12 years to maturity, a par value of $1000 and a selling for $960 at this price the bonds yield 10% what must the coupon rate be on the bonds?

Ο 10.00% Ο 18.83% Ο 9,41% Ο 9.51%
Answer #1

3= c (lty) -1 + f 960= c (1.1) ²1 t 1000 (1012 ool (1)? 960 = c(6.81369) + 3180636818 > c= 960-3180630818 = 94.13 6.81369 = 9

Please do rate me and mention doubts in the comments section.